# Tracing Nitrogen Enrichment across Cosmic Time with JWST

### v2

@@ 5 Discussion
While our findings are certainly affected by observational biases and selection effects (discussed in Appendix D), we are in agreement with recent studies pointing to moderate nitrogen over-enhancement for the bulk galaxy population at high $z$ (Masters et al., 2014; Strom et al., 2017; Hayden-Pawson et al., 2022; Hayes et al., 2025). However, those results were based on strong-line measurements or stacks, and thus the present work – incorporating direct-method abundances from auroral lines from individual galaxies, alongside consistently calibrated strong-line diagnostics – provides a more robust confirmation of the underlying trend.
@@
Various scenarios have been proposed in the literature to explain elevated N/O ratios at low metallicity. For example, bursty star formation, where the temporal decoupling between oxygen production (dominated by massive stars on few Myr timescales) and nitrogen release (from intermediate-mass stars on $\sim$100 Myr timescales) can lead to transiently high N/O ratios (Köppen & Hensler, 2005; Belfiore et al., 2015; Bhattacharya & Kobayashi, 2025; McClymont et al., 2025). If star formation proceeds in short, intense bursts followed by quiescent phases, oxygen may be injected early while nitrogen continues to enrich the ISM after the burst has faded, temporarily boosting N/O. In this scenario, as pointed out by McClymont et al. (2025), N/O overabundance should correlate with a negative offset with respect to the MS.
@@
In the right panel of Figure 6, we show the N/O-O/H relation colour-coded by $\Delta MS$, which represents the offset of each galaxy SFR from the star formation main sequence (MS) prediction. Whenever stellar masses were provided in the original publications, we adopted those values; otherwise, we used the estimates from the DJA catalogue, obtained from SED fitting as described in Valentino et al. (2023) using eazy-py (Brammer et al., 2008). SFRs were estimated them from the Hα luminosity using the conversion of Kennicutt (1998), scaled to a Chabrier initial mass function (Chabrier, 2003). We adopt the MS relation from Popesso et al. (2023) at z=2.7, i.e. the median redshift of our sample. Most N-enhanced objects in our sample exhibit positive $\Delta MS$, rather than the suppressed star formation activity expected if the N/O enhancement were driven by a temporary quenching phase following a starburst. This trend may, however, be biased by observational selection effects, so we caution against over-interpreting it at this stage.
@@
A related scenario involves the selective removal of metals by Type II supernova–driven outflows. Previous studies (e.g. Sugimura et al., 2024; Pallottini et al., 2025) suggest that energetic supernova feedback can eject oxygen-rich material more efficiently than the nitrogen released later by LIMS. Simple chemical evolution models show that such differential outflows can explain N over-enhancement (Rizzuti et al., 2025). McClymont et al. (2025) argue that bursty star formation naturally produced differential outflows when integrated over time, because the strongest feedback phase acts on the ISM after oxygen is newly synthetized. The gas-poor state following the burst can then be enriched in nitrogen by LIMS.
@@
A second class of explanations involves changes in gas metallicity driven by the inflow of pristine material. The accretion of metal-poor gas dilutes the ISM oxygen abundance, while leaving N/O essentially unaffected. This mechanism is expected to be particularly relevant at high redshift, when cosmological accretion rates are higher and cold flows are predicted to be common (Dekel et al., 2009; Tacconi et al., 2010; Davé et al., 2012). In this picture, galaxies experiencing substantial inflows should exhibit lower O/H than typical systems of the same stellar mass, while their N/O ratios should reflect the pre-inflow chemical state.
@@
In the left panel of Figure 6 we examined the N/O–O/H relation colour-coded by stellar mass. If dilution by pristine gas were the dominant mechanism, the most N-enhanced systems should also be the most massive: a pure inflow event primarily reduces O/H, shifting galaxies horizontally toward lower metallicity at fixed mass. In this scenario, a galaxy would retain the mass and N/O expected for an evolved, enriched system, while appearing more metal-poor. We quantified the stellar mass dependence across the N/O-O/H plane by calculating the gradient angle [^3] The gradient angle and significance were derived by fitting the directional derivative of M in the 2D plane and estimating uncertainties through bootstrap reshuffling of the mass values., obtaining a value of $12.7^{\circ}\pm 6.3^{\circ}$. The associated statistical significance ($p_{\text{value}}<0.05$) confirms the robustness of this trend. This trend means that N-enhanced galaxies tend to be preferentially associated with lower stellar masses, in direct contradiction of the dilution hypothesis. The current statistics are limited, and a larger sample will be needed to confirm this behaviour robustly. Nevertheless, the observed trend disfavours pristine-gas inflow as the primary driver of the nitrogen enhancement (see also Hayden-Pawson et al., 2022).
@@
Another relevant class of models invokes intermittent, non-monotonic star-formation histories coupled with galactic winds. For example, a dual-burst implementation was developed by Kobayashi & Ferrara (2024) to explain the extremely N-enhanced system GN-z11 at $z=10.6$ (Bunker et al., 2023; Cameron et al., 2023), where early nitrogen excess is attributed to Wolf–Rayet (WR) enrichment (see also Curti et al., 2025c, for an application of such a model to a $z=9.4$ galaxy). Such a temporary nitrogen enhancement scenario (probing timescales of less than about 100 Myr), however, requires specific timing: WR stars emerge only within a narrow evolutionary phase shortly after a starburst, and the resulting N enrichment is short-lived (Kobayashi & Ferrara, 2024). Although such coincidences may occur in individual systems, it is unlikely that this mechanism could operate in a coordinated way across a large galaxy sample, making it implausible as the primary driver of the systematic N/O elevation we observe.
@@
A similar dual-burst framework was applied by Bhattacharya & Kobayashi (2025) to reproduce the abundance patterns of low-redshift ( ¡ $z$ ¿ = 0.1), nitrogen-enriched, low-metallicity galaxies observed in DESI. We note that, contrary to Kobayashi & Ferrara (2024), in this case the observed N/O enhancement is a consequence of the combined effect of delayed N-enrichment from AGB stars and strong outflows occurring after the second burst of star-formation (which follows a gas accretion episode that significantly diluted the metallicity of the system), while the fast N-enhancement from WR stars444which happens before the O production from CCSNe, although included in the model, does not produce significant observable features due to the very short timescales involved.
@@
Figure 7: Same as Figure 4, but with Milky Way globular cluster contours (representing probability levels of 5, 16, 50, 84, and 95 percent), respectively of the first (in orange) and second (in green) generation of stars. The contours are taken from Ji et al. (2025), extracted from the APOGEE catalogue in SDSS DR17. The blue, dashed rectangle marks the region in which the majority of the sample of $z>4$ N/O-enhanced galaxies with abundance from UV emission lines from Ji et al. (2025) are located.
@@
Finally, several studies have proposed a link between nitrogen-enhanced high-redshift galaxies and the progenitors of present-day globular clusters (GCs) (e.g., Senchyna et al., 2023; Charbonnel et al., 2023; D’Antona et al., 2025). The unusual abundance patterns in some early galaxies resemble those in Milky Way GCs, suggesting that similarly dense, compact stellar systems might have influenced their chemical evolution. This connection was strengthened by Ji et al. (2025), who showed that the most N-enhanced $z>4$ galaxies whose abundances are measured with UV nitrogen lines share chemical signatures with second-generation (2G) GC stars. As shown in Figure 7, our comparison with Galactic GC populations reveals that our high-z ($z\sim 2$–4) sample, particularly direct-measurement galaxies, overlaps well with first-generation (1G) GC stars in the N/O–O/H plane, contrasting with the most metal-poor, ultra–N-enhanced $z>4$ systems, which overlap with 2G stars (Ji et al., 2025). This highlights that our objects display only moderate enhancement, incompatible with the extreme 2G-like signatures in the most UV-bright high-$z$ outliers. In this context, recent work suggests that 1G GC star formation may have occurred later than 2G (Clontz et al., 2024; McClymont et al., 2025), making a resemblance between 1G-like abundances and lower-redshift galaxies plausible. The physical origin of extreme nitrogen enrichment in 2G stars remains debated (e.g., Charbonnel et al., 2023), with scenarios such as super-massive star contributions still under discussion. This comparison should thus be regarded as phenomenological rather than evidence for a specific mechanism.
@@ 6 Conclusions
We presented a comprehensive investigation of the nitrogen enrichment in $\sim 500$ star-forming galaxies at redshift $z\approx 2-6$, combining ultra-deep JWST/NIRSpec spectroscopy from the MARTA survey with a large compilation of both high-$z$ and local Universe literature data. By directly detecting faint auroral lines in galaxies, we determined gas-phase abundances through the direct electron-temperature method, and recalibrated empirical strong-line relations for N/O at high redshift for the first time. Our work thus provides a homogeneous framework for comparing nitrogen enrichment across cosmic time, bridging the local Universe and galaxies up to $z\sim 8$. In particular, our main results include:
@@
Sample and methodology. The analysis combines 76 high-$z$ galaxies with auroral-line detections and 430 additional high-$z$ systems with strong-line measurements, together with $\sim 9\times 10^{4}$ local galaxies for comparison (Sect.2; Fig. 1). Direct Te-based abundances were derived when possible, while strong-line calibrations were applied elsewhere, following the self-consistent framework summarised in Fig. 2.
@@
Calibration of N/O strong-line diagnostics (Sect. 4.1; Fig. 3). We recalibrated the N2O2 and N2S2 diagnostics at high redshift using galaxies with direct N/O measurements. No systematic offset is observed relative to local calibrations, confirming that strong-line N/O indicators remain valid up to $z\sim 6$. Among the two diagnostics, N2O2 provides the tightest correlation, with an intrinsic scatter of $\sigma_{\rm int}\approx 0.09$ dex, while N2S2 is slightly less tight ($\sigma_{\rm int}\approx 0.13$ dex) but has the advantage of being less sensitive to dust extinction effects.
@@
The N/O-O/H relation across cosmic time (Sect. 4.2; Fig. 4, 5). Combining all datasets, we constructed the most extensive N/O-O/H relation for galaxies at $z>1$; we fitted the functional form from Hayden-Pawson et al. (2022) (Equation 5) to the local Universe dataset, finding the best-fit parameters reported in Table 1. High-redshift galaxies display a systematic enhancement of N/O relative to the local relation, by $\Delta(N/O)\approx 0.18$ dex (median), confirmed at >99.9>99.9% confidence (AD and Mann–Whitney tests). The enhancement is metallicity-dependent, reaching $\sim$0.3–0.4 dex at $12+\log(\mathrm{O/H})\lesssim 8.1$ (Fig. 5).
@@
Physical interpretation of N/O enhancement (Sect. 5; Fig. 6, 7). We explored several physical mechanisms proposed to explain the elevated N/O ratios at high redshift. Most N-enhanced galaxies in our sample lie above the main sequence (Fig. 6, right panel), which appears inconsistent with the expectation that bursty star formation followed by quiescent phases drives the N/O enhancement, since such scenario would predict suppressed star formation activity. We also find that N-enhanced systems are preferentially associated with lower stellar masses (Fig. 6, left panel), disfavouring pristine gas dilution as the main driver of the enhancement. Finally, comparison with Galactic globular cluster stellar populations (Fig. 7) shows that our $z\sim 2$–4 systems occupy the same region of the N/O–O/H diagram as first-generation GC stars, and remain well separated from the locus of 2G-like, ultra–N-enhanced high-redshift galaxies. This indicates that the moderate N/O excess we observe is consistent with relatively mild processing in compact star-forming regions, without invoking the extreme enrichment associated with the most chemically exceptional early-universe systems.
@@
Overall, this study establishes a comprehensive view of nitrogen enrichment at Cosmic Noon (<z>=2.7<z>=2.7), revealing the the overall population is mildly nitrogen enhanced, especially at low metallicity. Larger samples are needed to reduce observational bias, which may be driving some of the population properties observed here. However, taken at face value, none of the scenarios outlined in the literature fit our data in detail, outlining the need for renewed theoretical efforts in the field of chemical evolution modelling at Cosmic Noon and not only at $z>4$. The measurement of abundances of additional species will provide even more stringent constrains on theoretical work and greatly contribute to our understanding of how star formation proceeded from Cosmic Dawn to Noon.

### v3

@@ 5 Discussion
While our findings are certainly affected by observational biases and selection effects (discussed in Appendix D), we are in agreement with recent studies pointing to moderate nitrogen over-enhancement for the bulk galaxy population at high $z$ (Masters et al., 2014; Strom et al., 2017; Hayden-Pawson et al., 2022; Hayes et al., 2025; Cameron et al., 2026). However, those results were based on strong-line measurements or stacks, and thus the present work – incorporating direct-method abundances from auroral lines from individual galaxies, alongside consistently calibrated strong-line diagnostics – provides a more robust confirmation of the underlying trend.
@@
Various scenarios have been proposed in the literature to explain elevated N/O ratios at low metallicity. For example, bursty star formation, where the temporal decoupling between oxygen production (dominated by massive stars on few Myr timescales) and nitrogen release (from intermediate-mass stars on $\sim$100 Myr timescales) can lead to transiently high N/O ratios (Köppen & Hensler, 2005; Belfiore et al., 2015; Bhattacharya & Kobayashi, 2026; McClymont et al., 2026). If star formation proceeds in short, intense bursts followed by quiescent phases, oxygen may be injected early while nitrogen continues to enrich the ISM after the burst has faded, temporarily boosting N/O. In this scenario, as pointed out by McClymont et al. (2026), N/O overabundance should correlate with a negative offset with respect to the MS.
@@
In the left panel of Figure 6, we show the N/O-O/H relation colour-coded by $\Delta MS$, which represents the offset of each galaxy SFR from the star formation main sequence (MS) prediction. Whenever stellar masses were provided in the original publications, we adopted those values; otherwise, we used the estimates from the DJA catalogue, obtained from SED fitting as described in Valentino et al. (2023) using eazy-py (Brammer et al., 2008). SFRs were estimated then from the luminosity using the conversion of Kennicutt (1998), scaled to a Chabrier initial mass function (Chabrier, 2003). We adopt the MS relation from Popesso et al. (2023), evaluating it at the redshift of each individual galaxy in order to compute the corresponding $\Delta MS$. A high fraction of N-enhanced objects in our sample exhibit positive $\Delta MS$, rather than the suppressed star formation activity expected if the N/O enhancement were driven by a temporary quenching phase following a starburst. This trend may, however, be biased by observational selection effects, so we caution against over-interpreting it at this stage. It is also worth noting that, although the majority of galaxies ($\gtrsim$50%) with enhanced N/O lie above the main sequence, a non-negligible fraction is found below it, and overall we do not observe any clear correlation between $\Delta\log(\mathrm{N/O})$ and $\Delta MS$, as discussed in greater detail in Appendix D.
@@
A related scenario involves the selective removal of metals by Type II supernova-driven outflows. Previous studies (e.g. Sugimura et al., 2024; Pallottini et al., 2025) suggest that energetic supernova feedback can eject oxygen-rich material more efficiently than the nitrogen released later by LIMS. Simple chemical evolution models show that such differential outflows can explain N over-enhancement (Rizzuti et al., 2025). McClymont et al. (2026) argue that bursty star formation naturally produced differential outflows when integrated over time, because the strongest feedback phase acts on the ISM after oxygen is newly synthetized. The gas-poor state following the burst can then be enriched in nitrogen by LIMS.
@@
A second class of explanations involves changes in gas metallicity driven by the inflow of pristine material. The accretion of metal-poor gas dilutes the ISM oxygen abundance, while leaving N/O essentially unaffected. This mechanism is expected to be particularly relevant at high redshift, when cosmological accretion rates are higher and cold flows are predicted to be common (Dekel et al., 2009; Tacconi et al., 2010; Davé et al., 2012). In this picture, galaxies experiencing substantial inflows should exhibit lower O/H than typical systems of the same stellar mass, while their N/O ratios should reflect the pre-inflow chemical state. In particular, if the pre-inflow metallicity lies in the regime where secondary nitrogen production dominates, a galaxy can retain a high N/O ratio even after its O/H is temporarily diluted by a massive inflow of metal-poor gas.
@@
In the right panel of Figure 6 we examined the N/O–O/H relation colour-coded by stellar mass. If dilution by pristine gas were the dominant mechanism, the most N-enhanced systems should also be, generally, more massive: a pure inflow event primarily reduces O/H, shifting galaxies horizontally toward lower metallicity at fixed mass. In this scenario, a galaxy would retain the mass and N/O expected for an evolved, enriched system, while appearing more metal-poor. Conversely, we observe that lower-mass galaxies tend to have lower metallicities and vice versa, broadly consistent with the expectations from the mass-metallicity relation, without obvious systematic deviations that would suggest pristine gas dilution as the primary driver. We quantified the stellar mass dependence across the N/O-O/H plane by calculating the gradient angle [^4] The gradient angle and significance were derived by fitting the directional derivative of M in the 2D plane and estimating uncertainties through bootstrap reshuffling of the mass values., obtaining a value of $12.7^{\circ}\pm 6.3^{\circ}$. The associated statistical significance ($p_{\text{value}}\lesssim 10^{-6}$) confirms the robustness of this trend. This trend confirms that N-enhanced galaxies tend to be preferentially associated with lower stellar masses, which runs opposite to the expectations under an idealized pristine gas dilution scenario. The current statistics are limited, and a larger sample will be needed to confirm this behaviour robustly. Nevertheless, the observed trend disfavours pristine-gas inflow as the primary driver of the nitrogen enhancement (see also Hayden-Pawson et al., 2022).
@@
Another relevant class of models invokes intermittent, non-monotonic star-formation histories coupled with galactic winds. For example, a dual-burst implementation was developed by Kobayashi & Ferrara (2024) to explain the extremely N-enhanced system GN-z11 at $z=10.6$ (Bunker et al., 2023; Cameron et al., 2023), where early nitrogen excess is attributed to Wolf–Rayet (WR) enrichment (see also Curti et al., 2025c, for an application of such a model to a $z=9.4$ galaxy). Furthermore, direct observational signatures of WR stars – such as broad blue and orange emission bumps – have also been detected in several N-enhanced systems (e.g., Rivera-Thorsen et al., 2024; Berg et al., 2025; Morishita et al., 2025). However, such a temporary nitrogen enhancement scenario requires specific timing: WR stars, which are massive stars ($\gtrsim 25M_{\odot}$) in a very short-lived evolutionary phase, emerge only within a few Myr after a starburst and enrich the ISM on similarly rapid timescales (e.g., Schaerer & Vacca 1998; Crowther 2007). Although such coincidences may occur in individual systems, it is unlikely that this mechanism could operate in a coordinated way across a large galaxy sample, making it implausible as the primary driver of the systematic N/O elevation we observe.
@@
A similar dual-burst framework was applied by Bhattacharya & Kobayashi (2026) to reproduce the abundance patterns of low-redshift ( ¡ $z$ ¿ = 0.1), nitrogen-enriched, low-metallicity galaxies observed in DESI. We note that, contrary to Kobayashi & Ferrara (2024), in this case the observed N/O enhancement is a consequence of the combined effect of delayed N-enrichment from AGB stars and strong outflows occurring after the second burst of star-formation (which follows a gas accretion episode that significantly diluted the metallicity of the system), while the fast N-enhancement from WR stars555which happens before the O production from CCSNe, although included in the model, does not produce significant observable features due to the very short timescales involved.
@@
Figure 7: Same as Figure 4, but with Milky Way globular cluster contours (representing probability levels of 5, 16, 50, 84, and 95 percent), respectively of the first (in green) and second (in red) generation of stars. The contours are taken from Ji et al. (2026), extracted from the APOGEE catalogue in SDSS DR17.
@@
Finally, several studies have proposed a link between nitrogen-enhanced high-redshift galaxies and the progenitors of present-day globular clusters (GCs) (e.g., Charbonnel et al., 2023; Senchyna et al., 2024; D’Antona et al., 2025). The unusual abundance patterns in some early galaxies resemble those in Milky Way GCs, suggesting that similarly dense, compact stellar systems might have influenced their chemical evolution. This connection was strengthened by Ji et al. (2026), who showed that the most N-enhanced $z>4$ galaxies whose abundances are measured with UV nitrogen lines share chemical signatures with second-generation (2G) GC stars. As shown in Figure 7, our comparison with Galactic GC populations reveals that our high-z ($z\sim 2$–4) sample, particularly direct-measurement galaxies, overlaps better with first-generation (1G) GC stars in the N/O–O/H plane, contrasting with the most metal-poor, ultra–N-enhanced $z>4$ systems, which overlap with 2G stars (Ji et al., 2026), highlighting that our objects display only moderate enhancement, incompatible with the extreme 2G-like signatures in the most UV-bright high-$z$ outliers. We note, however, that the median properties of 1G GC stars are broadly consistent with the local N/O-O/H relation. Our galaxies therefore fall within the intrinsic scatter of the 1G distribution rather than tracing its median trend. A simple quantitative test shows that $\sim$70% of the high-z objects lie within the 84th percentile of the 1G distribution, although their median properties remain offset.
In this context, recent work suggests that 1G GC star formation may have occurred later than 2G (Clontz et al., 2024; McClymont et al., 2025), making a resemblance between 1G-like abundances and lower-redshift galaxies plausible. The physical origin of extreme nitrogen enrichment in 2G stars remains debated (e.g., Charbonnel et al., 2023), with scenarios such as super-massive star contributions still under discussion. This comparison should thus be regarded as phenomenological rather than evidence for a specific mechanism.
@@ 6 Conclusions
We presented a comprehensive investigation of the nitrogen enrichment in $\sim 500$ star-forming galaxies at redshift $z\approx 2-4$, combining ultra-deep JWST/NIRSpec spectroscopy from the MARTA survey with a large compilation of both high-$z$ and local Universe literature data. By directly detecting faint auroral lines in galaxies, we determined gas-phase abundances through the direct electron-temperature method, and recalibrated empirical strong-line relations for N/O at high redshift for the first time. Our work thus provides a homogeneous framework for comparing nitrogen enrichment across cosmic time, bridging the local Universe and galaxies up to $z\sim 8$. In particular, our main results include:
@@
Sample and methodology. The analysis combines 76 high-$z$ galaxies with auroral-line detections and 430 additional high-$z$ systems with strong-line measurements, together with $\sim 9\times 10^{4}$ local galaxies for comparison (Sect.2; Figure 1). Direct Te-based abundances were derived when possible, while strong-line calibrations were applied elsewhere, following the self-consistent framework summarised in Figure 2.
@@
Calibration of N/O strong-line diagnostics (Sect. 4.1; Figure 3). We recalibrated the N2O2 and N2S2 diagnostics at high redshift using galaxies with direct N/O measurements. No systematic offset is observed relative to local calibrations, confirming that strong-line N/O indicators remain valid up to $z\sim 6$. Among the two diagnostics, N2O2 provides the tightest correlation, with an intrinsic scatter of $\sigma_{\rm int}\approx 0.09$ dex, while N2S2 is slightly less tight ($\sigma_{\rm int}\approx 0.13$ dex) but has the advantage of being less sensitive to dust extinction effects.
@@
The N/O-O/H relation across cosmic time (Sect. 4.2; Figure 4, 5). Combining all datasets, we constructed the most extensive N/O-O/H relation for galaxies at $z>1$; we fitted the functional form from Hayden-Pawson et al. (2022) (Equation 5) to the local Universe dataset, finding the best-fit parameters reported in Table 1. High-redshift galaxies display a systematic enhancement of N/O relative to the local relation, by $\Delta(N/O)\approx 0.18$ dex (median), confirmed at >99.9>99.9% confidence (AD and Mann–Whitney tests). The enhancement is metallicity-dependent, reaching $\sim$0.3–0.4 dex at $12+\log(\mathrm{O/H})\lesssim 8.1$ (Figure 5).
@@
Physical interpretation of N/O enhancement (Sect. 5; Figure 6, 7). We explored several physical mechanisms proposed to explain the elevated N/O ratios at high redshift. Most N-enhanced galaxies in our sample lie above the main sequence (Figure 6, right panel), which appears inconsistent with the expectation that bursty star formation followed by quiescent phases drives the N/O enhancement, since such scenario would predict suppressed star formation activity. We also find that N-enhanced systems are preferentially associated with lower stellar masses (Figure 6, left panel), disfavouring pristine gas dilution as the main driver of the enhancement. Finally, comparison with Galactic globular cluster stellar populations (Figure 7) shows that our $z\sim 2$–4 systems lie within the scatter of first-generation GC stars, yet their median N/O at a given metallicity exceeds the 1G median trend, and remain well separated from the locus of 2G-like, ultra–N-enhanced high-redshift galaxies. This indicates that the moderate N/O excess we observe is consistent with relatively mild processing in compact star-forming regions, without invoking the extreme enrichment associated with the most chemically exceptional early-universe systems.
@@
Overall, this study establishes a comprehensive view of nitrogen enrichment at Cosmic Noon (<z>=3.0<z>=3.0), revealing the the overall population is mildly nitrogen enhanced, especially at low metallicity. Larger samples are needed to reduce observational bias, which may be driving some of the population properties observed here. However, taken at face value, none of the scenarios outlined in the literature fit our data in detail, outlining the need for renewed theoretical efforts in the field of chemical evolution modelling at Cosmic Noon and not only at $z>4$. The measurement of abundances of additional species will provide even more stringent constrains on theoretical work and greatly contribute to our understanding of how star formation proceeded from Cosmic Dawn to Noon.